### v1

# Identifying Coleman-Weinberg type-chiral phase transition in QCD-like scenario

Abstract
The chiral phase transition in QCD-like theories can be Coleman-Weinberg type supercooling. This is possible to see when Beyond the Standard Model (say, axionlike particle or baryogenesis) is coupled to QCD of the Standard Model (called ordinary QCD). Identifying such a supecooling chiral phase transition thus opens a new window to search for new physics associated with the QCD phase transition epoch of the thermal history, to be probed by gravitational wave and primordial black hole productions. We discuss this possibility by monitoring ordinary QCD setup in a view of a two-flavor Nambu-Jona-Lasinio model in hot and/or dense medium, at the mean-field approximation level. We find that the identification essentially follows the scale violation classification: soft-scale breaking and quantum scale anomaly.
I Introduction
First order phase transitions or supercooling phase transitions currently involve broader interests over theoretical particle physics and cosmology Mazumdar:2018dfl . A particular motivation can be seen in cosmological applications to the thermal history of the Universe: the first-order/supercooling phase transition serves as productions for stochastic gravitational wave (GW) background and primordial black hole (PBH) formation Jung:2021mku ; Kodama:1982sf ; Hall:1989hr ; Kusenko:2020pcg ; Hashino:2021qoq ; Hashino:2022tcs ; He:2022amv ; Liu:2021svg ; Kawana:2022olo ; Lewicki:2023ioy ; Gouttenoire:2023naa ; Gouttenoire:2023pxh ; Salvio:2023blb ; Conaci:2024tlc ; Banerjee:2023brn ; Banerjee:2023vst ; Banerjee:2023qya ; Lewicki:2024ghw ; Flores:2024lng ; Cai:2024nln ; Florentino:2024kkf ; Goncalves:2024vkj ; Banerjee:2024fam ; Shao:2024dxt ; Banerjee:2024cwv ; Wu:2024lrp ; Hashino:2025fse ; Murai:2025hse ; Franciolini:2025ztf ; Kierkla:2025vwp ; Kanemura:2024pae as nonparticle dark matter candidate Carr:2020gox ; Baldes:2023rqv ; Balaji:2024rvo ; Jinno:2023vnr ; Zhang:2025kbu ; Kanemura:2024pae . It has also been argued that the supercooling electroweak or QCD phase transition can serve as a possible initial condition for slow-roll inflation Zhang:2023acu ; Liu:2024xrh , and subsequent particle production Wang:2022ygk ; Wang:2024tda .
One typical modeling to access those cosmological applications would be a scale-symmetry phase transition via the Coleman-Weinberg (CW) mechanism Coleman:1973jx , assuming an appropriate presence of scalar field (presumably allowing coupled to the standard model plasma). All dimensionful parameters are absent at tree level, but quantum loop corrections induce the scale anomaly and generate the CW potential leading to the spontaneous scale-symmetry breaking. Since the CW potential merely includes a logarithmic field dependence, it exponentially suppresses the tunneling rate at low temperatures. This intrinsic feature naturally induces strong enough supercooling, hence is particularly relevant for the PBH formation Gouttenoire:2023naa ; Gouttenoire:2023pxh ; Baldes:2023rqv ; Conaci:2024tlc ; Arteaga:2024vde ; Salvio:2023ynn ; Salvio:2023blb ; Banerjee:2024cwv ; Kierkla:2025vwp .
Very recently, the importance of what is called the “soft scale-breaking” mass term has also been emphasized in addressing the CW-type supercooling phase transition Zhang:2025kbu . This soft-scale breaking term can dramatically alter the prediction for PBH formation, due to violation of the exponential nucleation approximation. However, it still leaves the successful stochastic GW background predictions essentially as they stand in the case of the pure CW scenario Jaeckel:2016jlh ; Marzola:2017jzl ; Iso:2017uuu ; Ghorbani:2017lyk ; Baldes:2018emh ; Prokopec:2018tnq ; Brdar:2018num ; Marzo:2018nov ; DelleRose:2019pgi ; Ghoshal:2020vud ; Levi:2022bzt ; Kierkla:2022odc ; Salvio:2023qgb ; Zhang:2024vpp . No matter how small such a deformation is, the soft-scale breaking is still active even at low enough temperature – much lower than the critical temperature for the thermal phase transition ($T_{c}$), but as low as the critical temperature for the bubble nucleation ($T_{n}$). Thus this deformation possibility has provided a new testable link between PBH dark matter and GW signatures in the CW-type scenario.
The potential form of the deformed CW potential addressed in Zhang:2025kbu essentially goes like
$V(M:T\ll T_{c})\sim m_{0}^{2}M^{2}+c_{1}\cdot M^{4}[\ln(M/v)+c_{2}]\,.$ (1)
Here $M$ is a CW scalar field; the $m_{0}^{2}$ term corresponds to what is called the soft-scale breaking term; $v$ denotes the vacuum expectation value of $M$; $c_{1,2}$ are parameters to be determined at one-loop level given a model. The potential barrier around the origin can be generated when $m_{0}^{2}>0$ with $c_{1}>0$, much like the case of the CW thermal phase transition. As long as $m_{0}^{2}$ is small enough, this phase transition will be supercooled due to the approximate scale invariance around the origin of the potential.
Actually, the potential form Eq.(1) is also the one that arises in low-energy chiral effective models of QCD and QCD-like theories, such as a Nambu-Jona-Lasinio (NJL) description in the mean-field approximation (MFA) with the bosonization of chiral quark bilinear: $M\sim\bar{q}q$. In there, the logarithmic term $\sim M^{4}\ln M$ is generated as a consequence of the quantum scale anomaly, which is also related to the renormalization group evolution of the dynamically induced quartic coupling for $M$. In medium with, e.g., finite baryon chemical potential $\mu_{B}$, this term would get corrections depending on $\mu_{B}$. The soft-scale breaking term $\sim m_{0}^{2}M^{2}$, at $T\sim T_{n}\ll T_{c}$, would be governed by the vacuum contribution, which induces $m_{0}^{2}<0$ as the dynamical chiral symmetry breaking, and also Debye screening mass in medium, $\sim\mu_{B}^{2}M^{2}$, which drives $m_{0}^{2}$ to be positive. Furthermore, beyond the Standard Model contribution, if coupled to the QCD chiral phase transition, might also come in the soft-scale breaking term. Thus, the chiral phase transition in QCD-like theories can also be argued in terms of the same class of the scale violation, the soft-scale breaking $\sim M^{2}$ and the quantum scale anomaly $\sim M^{4}\ln M$, as discussed in Zhang:2025kbu .
In QCD of the Standard Model, which we shall call ordinary QCD, the chiral phase transition goes like crossover at around vanishing baryon chemical potential Aoki:2006we ; Borsanyi:2020fev ; Steinbrecher:2018phh . Thereby, a first-order and supercooling phase transition would be unlikely to take place within ordinary QCD framework. This feature is also reflected in the MFA-NJL description (see, e.g., reviews Hatsuda:1994pi ). The deviation from the CW-type phase transition is tightly linked with the fact that ordinary QCD generates a negatively too large soft-scale breaking mass ($m_{0}^{2}<0$ in Eq.(1)), as will be more explicitized later.
In the literature Shao:2024dxt ; Shao:2024ygm , the possibility of realizing the supercooling from the QCD phase transition has been explored based on a couple of low-energy effective models. There it has been clarified that with a baryon chemical potential much higher than the typical QCD scale $(\lesssim 200-400\,{\rm MeV})$, a small phase transition rate $(\beta/H)$ enough to produce nano hertz GWs can be achieved. This, in terms of the soft-scale breaking, occurs due to not merely the creation of the Fermi surface, but the Debye mass $\sim\mu_{B}^{2}M^{2}$ that is positive and still large enough even at $T\sim T_{n}$, and also.
Another Debye mass effect as a positive soft-scale breaking has also been discussed in the recent literature Zhang:2025kbu . In the reference, in a context different from ordinary QCD, the chiral (axial) chemical potential $\mu_{5}$ in a model of scale-invariant dark QCD does the job along with mixing with an axial charged axionlike particle (ALP).
Thus it would be suspected that the scale violation classification, as briefly illustrated above, can give a unified identifier of the chiral phase transition and a discriminator for the first-order. More remarkably, one could identify the CW-type chiral phase transition even in QCD-like theories. This might be required to have new physics coupled to QCD or QCD-like theories. Pursing this possibility would pave a new way toward hints for new physics coupled or related to the QCD chiral phase transition in light of GW productions and PBH formation.
In this paper, we attempt to clarify how the chiral phase transition in QCD-like theories can be CW-type with Beyond the Standard Model. Our investigation urges us to propose three scenarios: i) many-flavor (walking) dark QCD; ii) QCD with an ALP; iii) the chiral phase transition with a large $\mu_{B}$ $\sim 1$ GeV, where, as done in Shao:2024dxt ; Shao:2024ygm , the Affeck-Dine baryogenesis Affleck:1984fy ; Linde:1985gh ; Dine:2003ax is assumed to leave a large amount of $\mu_{B}$ in the QCD phase transition epoch. In all those cases, the scale violation classification works well as well as the CW-type scenario.
The CW-type supercooling phase transition can be realized even in QCD-like theories, where the phase transition essentially takes the potential form in Eq.(1). This is possible as long as the quantum scale anomaly takes the four- or three-dimensional form. The strongly magnetized medium is, therefore, still nontrivial because the four-dimensional system is to be reduced to the two-dimensional one characterized as $SO(2)\times SO(1,1)$ symmetry. This case will be disregarded in the present study.
We employ a two-flavor NJL model in the MFA, with monitoring the QCD case, and analyzing the thermal chiral phase transition in the presence of the chiral chemical potential $\mu_{5}$, nonzero strong CP phase $\theta$, in addition to the baryon chemical potential $\mu_{B}$. The scale violation classification is also operative in medium characterized by other chemical potentials, such as the isospin chemical potential ($\mu_{I}$), as long as the extreme condition is supplied via external gauging of the global chiral symmetry.
The three-flavor QCD case would have another driving force of the trilinear term $M^{3}$, which can also generate a potential barrier, associated with the $U(1)$ axial anomaly and the QCD instanton Kobayashi:1970ji ; Kobayashi:1971qz ; tHooft:1976snw . This cubic term, however, is not essential for the spontaneous breaking of the chiral symmetry – which does not destabilize the symmetric phase at $M=0$ and acts as higher order correction, when the negative $M^{2}$ term is present – and does not get thermal corrections from quark loop at the MFA level. Therefore, this cubic term would be subdominant when the CW-type chiral phase transition is realized and the supercooling is processed.
Here we highlight generic features that will be clarified in the present paper:
  • In the NJL-like models, the several types of sources for “soft” scale violation, up to temperature $T$, are introduced as
    i) charge or matter density via gauging global currents (in part); e.g. $\mu_{B}$, $\mu_{5}$, the Polyakov loop (as an imaginary chemical potential),
    ii) current quark mass terms $m$ and $me^{i\theta}$;
    iii) dynamical chiral symmetry breaking via four-fermion interactions.
  • At high enough $T$, the thermal correction cancels the CW-type scale anomaly of the form $\sim M^{4}\log M$, and the chiral phase transition is triggered simply by the generated Debye mass for $M$ and thus goes like the second-order type.
  • If the scale symmetry of a theory is only logarithmically broken, like CW-type models, the high-$T$ expansion is insufficient to see the thermal phase transition, which is because the transition can take place at relatively lower $T$ in comparison with the case where a soft scale-breaking, negative mass-squared term $(\sim-M^{2})$ is present. In that case, the logarithmic term is not perfectly canceled enough to interfere creation of the potential barrier around $M=0$. Thus this type of theory undergoes a first-order phase transition, just like the CW-type models.
  • In comparison with the CW-type model, QCD-like theories induce the dynamical chiral symmetry breaking, which reflects the negative mass-squared term for $M$ in the MFA-NJL description or bosonization. This deepens the curvature around $M=0$, so that the transition temperature becomes high enough that the high-$T$ expansion works well and cancels the net logarithmic term to wash out the barrier wall. Thus this causes the second-order, or crossover phase transition. This is why the first order is not realized along the $T$-axis in QCD-like theories. This also implies a key understanding of how the electroweak phase transition becomes crossover in the Standard Model with the negative soft-scale breaking term $\sim-H^{2}$.
  • Going along the $\mu_{B}$ axis at $T=0$, the CW-type scale anomaly of the form $\sim M^{4}\log M$ is also canceled by the finite density ($\mu_{B}$) correction, in a way similar to the high $T$ correction case. Still, the finite $\mu_{B}$ generates not only the Debye mass for $M$, but also a potential barrier due to the creation of the Fermi surface which acts as a repulsive force. Therefore, at $T=0$, the chiral phase transition is of first order which is not generically of the CW type form. However, at finite $T$, it is possible to have the CW-type phase transition with an appropriate $\mu_{B}$ by controlling the net soft-scale breaking mass term.
This paper is structured as follows: in Sec. II, we present the preliminary setup of the two-flavor NJL model and derive the thermodynamic potential in the MFA with nonzero $\mu_{5},\mu_{B},\theta$ at finite $T$. Section III clarifies the generic properties of the thermal chiral phase transition at $\mu_{B}=0$ in comparison with the CW-type phase transition, particularly paying our attention to the scale violation classes. Then, we propose two classes of “projection” from the chiral phase transition fom in ordinary QCD case into the CW supercooling form: walking dark QCD (Sec. III.1) and QCD with ALP (Sec. III.2). In Sec. IV, we discuss the case with finite $\mu_{B}$, and demonstrate that a large $\mu_{B}$ ($\sim 1$ GeV) can produce the Debye mass properly to make the chiral phase transition CW-typed. The conclusion of this paper is provided in Sec. V.
II Two-flavor mean-field NJL monitor
We employ a two-flavor NJL model with the lightest quark doublet $q=\left(u,d\right)^{T}$ and focus on the scale violation classes as has briefly been addressed in the Introduction. The Lagrangian is given as
$\mathcal{L}$ $=\bar{q}\left(i\gamma_{\mu}\partial^{\mu}-m+\frac{1}{3}\mu_{B}\gamma^{0}+\mu_{5}\gamma^{0}\gamma^{5}\right)q$
$+\frac{g_{s}}{2}\sum^{3}_{a=0}\left[\left(\bar{q}\tau_{a}q\right)^{2}+\left(\bar{q}i\gamma_{5}\tau_{a}q\right)^{2}\right]+g_{d}\left(e^{i\theta}\text{det}\left[\bar{q}\left(1+\gamma_{5}\right)q\right]+e^{-i\theta}\text{det}\left[\bar{q}\left(1-\gamma_{5}\right)q\right]\right)\,,$ (2)
where we have taken the isospin symmetric limit: $m_{u}=m_{d}=m$, and have also introduced the determinant term related to the $U(1)_{A}$ anomaly in the underlying QCD; $\tau_{a}$ denotes the Pauli matrices with $\tau_{0}=1_{2\times 2}$. The current quark mass is defined to be real and positive, so that the strong CP phase $\theta$ coupled to the $U(1)$ axial anomalous determinant term has already involved the quark mass phase, i.e., $\theta\equiv\bar{\theta}$. Regarding the scale violation classes at this point, as listed in the Introduction, we see that
  • i)
    the baryon chemical potential $\mu_{B}$, the chiral chemical potential $\mu_{5}$, arising via gauging the $U(1)$ vectorial and axial global symmetries, respectively;
  • ii)
    the current quark mass $m$;
  • iii)
    the four-fermion interactions along with the couplings $g_{s}$ and $g_{d}$, which are the essential driving force to trigger dynamical chiral symmetry breaking.
It is convenient to transform quark fields so as to remove $\theta$ from the $g_{d}$- determinant term, by a U(1) axial rotation $q\to q^{\prime}=e^{-i\gamma_{5}\frac{\theta}{4}}q$. The Lagrangian in Eq.(II) is then transformed as
$\mathcal{L}\to\mathcal{L}^{\prime}$ $=\bar{q}^{\prime}\left(i\gamma_{\mu}\partial^{\mu}-m\left[\text{cos}\frac{\theta}{2}+i\gamma_{5}\text{sin}\frac{\theta}{2}\right]+\frac{1}{3}\mu_{B}\gamma^{0}+\mu_{5}\gamma^{0}\gamma^{5}\right)q^{\prime}$
[…]
We work in the MFA and take into account the mean field only for the $SU(2)$ isospin singlet #1#1#1This assumption is ensured by absence of the isospin chemical potential. Our current discussions can also be straightforwardly extended by including the isospin breaking.. Then the targeted mean fields are those for the following bilinears: $\left(\bar{q}^{\prime}_{i}q^{\prime}_{i}\right)$ and $\left(\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\right)$. Those are expanded around the mean fields $S^{\prime}=\langle\bar{q}^{\prime}_{i}q^{\prime}_{i}\rangle$ and $P^{\prime}=\langle\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\rangle$ as $\bar{q}^{\prime}_{i}q^{\prime}_{i}=S^{\prime}+\left(:\bar{q}^{\prime}_{i}q^{\prime}_{i}:\right)$ and $\langle\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\rangle=P^{\prime}+\left(:\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}:\right)$2#2#2 The U(1) axial rotation relates the scalar and pseudoscalar bilinears between the original- and prime-base scalar and pseudoscalar bilinears (for each quark flavor $i$) as $\left(\bar{q}_{i}q_{i}\right)$ $=\left(\bar{q}^{\prime}_{i}q^{\prime}_{i}\right)\text{cos}\frac{\theta}{2}+\left(\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\right)\text{sin}\frac{\theta}{2}\,,$ $\left(\bar{q}_{i}i\gamma_{5}q_{i}\right)$ $=-\left(\bar{q}^{\prime}_{i}q^{\prime}_{i}\right)\text{sin}\frac{\theta}{2}+\left(\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\right)\text{cos}\frac{\theta}{2}\,.$ Here the terms sandwiched by ”:” stand for the normal order product, which means $\langle:\mathcal{O}:\rangle=0$ for $\mathcal{O}=S^{\prime},P^{\prime}$, hence ${\cal L}={\cal L}|_{\rm MF-only}+(:{\cal L}:)$. Thus, in the MFA, the Lagrangian in Eq.(II) takes the form
${\cal L}_{\rm MF-only}\equiv\mathcal{L}_{\text{MFA}}=\sum_{i}\bar{q}^{\prime}_{i}\left(i\gamma_{\mu}\partial^{\mu}-(\alpha+i\gamma_{5}\beta)+\frac{1}{3}\mu_{B}\gamma^{0}+\mu_{5}\gamma^{0}\gamma^{5}\right)q^{\prime}_{i}-2g_{s}\left(S^{\prime 2}+P^{\prime 2}\right)-2g_{d}\left(S^{\prime 2}-P^{\prime 2}\right)\,,$ (4)
where
$\alpha$ $=m\medspace\text{cos}\frac{\theta}{2}-2\left(g_{s}+g_{d}\right)S^{\prime}\,,$
$\beta$ $=m\medspace\text{sin}\frac{\theta}{2}-2\left(g_{s}-g_{d}\right)P^{\prime}\,.$ (5)
By integrating out quark fields and applying the imaginary time formalism, we get the thermodynamic potential $\Omega$,
$\Omega=2g_{s}\left(S^{\prime 2}+P^{\prime 2}\right)+2g_{d}\left(S^{\prime 2}-P^{\prime 2}\right)-N_{c}N_{f}\sum_{s=\pm 1}\int\frac{d^{3}p}{(2\pi)^{3}}\left[E_{s}+T\text{ln}\left(1+e^{-\frac{(E_{s}+\frac{1}{3}\mu_{B})}{T}}\right)+T\text{ln}\left(1+e^{-\frac{(E_{s}-\frac{1}{3}\mu_{B})}{T}}\right)\right]\,,$ (6)
where $N_{f}=2$, $N_{c}=3$ for the QCD monitor, and
$E_{s}=\sqrt{M^{2}+(|p|-s\mu_{5})^{2}}\,,\qquad M^{2}=\alpha^{2}+\beta^{2}\,.$ (7)
We stabilize this $\Omega$ via the stationary condition set for given $\theta$, $\mu_{B}$, $\mu_{5}$, and $T$: $\frac{\partial\Omega}{\partial S^{\prime}}=\frac{\partial\Omega}{\partial P^{\prime}}=0$.
As is evident in the form of Eq.(6), $\mu_{5}$ and $\theta$ merely act as a catalyzer against the CW-type scale anomaly, i.e., the diving force for the phase transition in the case without them. For instance, even for a large $\mu_{5}$ at $T=\mu_{B}=0$, in the case with massive quarks including physical point, $\mu_{5}$ does not generate the Debye mass or logarithmic term for $M$, hence $\mu_{5}$ itself does not trigger the chiral restoration. In the chiral limit, however, $\mu_{5}$ can act like $\mu_{B}$ as seen from Eq.(7), and therefore can trigger the first order phase transition as has been discussed in the literature Chernodub:2011fr . Even at physical point, a dynamically induced $\mu_{5}$ as the mean field of the axialvector field can also serve as the source for triggering the first order phase transition Shao:2022oqw , which we will not consider in the present study. In what follows, we will thus take $\theta$ and $\mu_{5}$ into account as an option.
III Exploring CW-type chiral phase transition: case with $\mu_{B}=0$
In this case Eq.(6) goes like
$\left.\Omega\right|_{\mu_{B}=\mu_{5}=0}=2g_{s}(S^{\prime 2}+P^{\prime 2})+2g_{d}(S^{\prime 2}-P^{\prime 2})-2N_{c}N_{f}\int\frac{d^{3}p}{(2\pi)^{3}}\left[E_{s=0}+2T\text{ln}(1+e^{-\frac{E_{s=0}}{T}})\right]\,.$ (8)
It is well known in the two-flavor NJL model that at the physical point for quark masses, the chiral phase transition at vanishing baryon chemical potentials is crossover (see, e.g., reviews Hatsuda:1994pi ). This is also true even with nonzero $\theta$, as long as $\theta\neq\pi$, at which case the CP phase transition along $P=S^{\prime}$ undergoes like second order Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs ; Sasaki:2013ewa ; Lu:2018ukl ; Huang:2024nbd . Now we understand this phase transition property in terms of the scale violation classes.
First of all, we take $\theta=0$ just for simplicity. It turns out that this will not essentially lose generality on the phase transition property even with $\theta\neq 0$, as long as $0<\theta<\pi$. With $\theta=0$, we have $S^{\prime}=S$ and $P^{\prime}=P=0$, or equivalently, ($\beta=0$ in Eq.(II), hence $M^{2}=\alpha^{2}$ in Eq.(7). In that case Eq.(8) can still further be simplified to
$\left.\Omega\right|_{\mu_{B}=\mu_{5}=0}^{\theta=0}=\frac{(M-m)^{2}}{2(g_{s}+g_{d})}-2N_{c}N_{f}\int\frac{d^{3}p}{(2\pi)^{3}}\left[E_{s=0}+2T\text{ln}(1+e^{-\frac{E_{s=0}}{T}})\right]\,.$ (9)
Now, this is simply a function of $M$.
Consider first the vacuum case with $T=0$. The tree-level, bare, $(M-m)^{2}$ term has come from the four-fermion interactions along with $g_{s}$ and $g_{d}$, including the shift by the current quark mass. This is a soft-scale breaking term. In addition, the first term in the loop correction part (inside the square bracket in Eq.(9)) includes quadratic $\sim M^{2}\Lambda^{2}$ and logarithmic divergent terms $\sim M^{4}\ln(\Lambda/M)$, where $\Lambda$ denotes the loop momentum cutoff. The former definitely takes negative sign, and gives a destructive contribution to the bare $M^{2}$ term, which destabilizes the chiral symmetric phase at $M=0$ up to the linear shift by $m$, hence triggers the dynamical chiral symmetry breaking. This is another soft-scale breaking term. The net soft-breaking term takes largely negative sign at $T=0$. Meanwhile, the latter logarithmic divergent term takes nothing but a form of the CW-type scale anomaly. Thus, at $T=0$, the effective potential derived from Eq.(9) (with renormalization prescribed) can be cast into essentially the same form as in the generic CW-type potential given in Eq.(1), with a large negative curvature at around $M=0$ ($m_{0}^{2}<0$ in Eq.(1)):
$\left.\Omega\right|_{\mu_{B}=\mu_{5}=0}^{\theta=0,T=0}\approx\frac{(M-m)^{2}}{2(g_{s}+g_{d})}-\frac{N_{f}N_{c}}{4\pi^{2}}\left(M^{2}\Lambda^{2}-\frac{M^{4}}{4}\ln\frac{\Lambda^{2}}{M^{2}}\right)\,,$ (10)
for large $\Lambda$.
At finite $T$, the negative curvature at round $M=0$ gets thermal corrections (the second term in square bracket in Eq.(9)). This starts to act as a cancellation of the soft-scale breaking term $\propto M^{2}$. At some high enough $T$, the original soft-breaking term, i.e. the dynamical chiral symmetry breaking will be canceled (i.e., $m_{0}^{2}=0$ in Eq.(1)), that determines the typical critical temperature $T_{c}$. As $T$ gets further higher, the soft-breaking term is instead developed by the thermal mass term $\sim T^{2}M^{2}$. At this moment, one might suspect that a potential barrier is built around $M=0$ by the positive curvature (the thermal mass) $\sim(+T^{2}M^{2})$ and the logarithmic damping term $\sim(-M^{4}\ln M)$, just like the case of the CW-type potential at high $T$, hence the thermal chiral phase transition goes like first order or more like a supercooling form. However, it is not the case because of the nontrivial cancellation for the original CW-type logarithmic term by the high $T$ correction.
To see this cancellation more explicitly, we work on the high $T$ expansion for the thermal correction term in Eq.(9), which goes like
$\left.-N_{c}N_{f}\cdot T\int\frac{d^{3}p}{(2\pi)^{3}}\ln(1+e^{-\frac{E_{0}}{T}})\right|_{T\gg M}$ $=N_{c}N_{f}\left[-\frac{7\pi^{2}}{180}T^{4}+\frac{1}{12}T^{2}M^{2}+\frac{1}{16\pi^{2}}M^{4}\ln M^{2}-\frac{1}{16\pi^{2}}M^{4}\ln a_{f}T^{2}+\cdots\right]\,,$ (11)
where $a_{f}=2\ln\pi-2\gamma_{E}\simeq 1.14$. As seen from Eqs.(10) and (11), the original CW-type logarithmic term $\sim M^{4}\ln M$ is precisely canceled each other. Thereby, the CW-type first-order phase transition is not realized: no barrier is generated. This is completely consistent with the well-known fact that the two-flavor chiral phase transition in the MFA-NJL model, with $\mu_{B}=0$, is second order in the chiral limit with $m=0$, and crossover at physical point. This phenomenon can also be understood in a view of the original scale anomaly cancellation in an extremely hot medium, which implies a free quark gas picture simply following the Stephan-Boltzmann form of pressure, $\Omega\sim p\sim T^{4}$ (corresponding to the first term in Eq.(11)).
This scale anomaly cancellation does not happen when the soft-scale breaking mass is small enough as in the case that has been addressed in the literature Gupta:2011ez . This is because the high $T$ expansion will not well approximate the thermal potential due to the scale invariance at around $M=0$, hence a barrier is created and the phase transition undergoes like first order and supercooling. Thus, the scale anomaly cancellation is intrinsic to a negatively large soft-scale breaking case, which makes the curvature of the potential deeply negative at around $M=0$, like in the present NJL scenario with the significant dynamical chiral symmetry breaking.
Even when nonzero $\theta$ is turned on, we can again arrange $\Omega$ so that the thermal loop and vacuum-one loop terms are still singly dependent of $M$. Nonzero $\theta$ effects only come into the tree-level part, in which either $P^{\prime}=\frac{m\sin\frac{\theta}{2}-\beta}{2(g_{s}-g_{d})}$ or $S^{\prime}=\frac{m\cos\frac{\theta}{2}-\alpha}{2(g_{s}+g_{d})}$ direction can be chosen as the actual mean field directions in addition to the $M=\sqrt{\alpha^{2}+\beta^{2}}$ direction. Thus, the dynamical chiral symmetry breaking is still monitored along the $M$ direction, with keeping the size of the soft-scale symmetry breaking at each $T$. Hence the high $T$ expansion works as in the case with $\theta=0$ and the scale anomaly cancellation is still operative, so the chiral phase transition substantially follows the same way and no first order phase transition can be seen even with $\theta\neq 0$. On the other hand, the $CP$ phase transition feature observed along $S^{\prime}$ or $\alpha$ highly depends on $\theta$, which exhibits a second order type when $\theta=\pi$ Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs ; Huang:2024nbd .
The dynamical chiral symmetry breaking occurs when the gap equation, i.e., the stationary condition for Eq.(10) gets nontrivial solution. It is the case when the following condition is met:
$\frac{N_{f}N_{c}(g_{s}+g_{d})\Lambda^{2}}{2\pi^{2}}\geq 1\,,$ (12)
in the chiral limit, for simplicity. At the criticality where $\frac{N_{f}N_{c}(g_{s}+g_{d})\Lambda^{2}}{2\pi^{2}}=1$, the soft-scale breaking mass becomes zero: $m_{0}^{2}=0$ in terms of Eq.(1). In two-flavor QCD case, we have $\frac{N_{f}N_{c}(g_{s}+g_{d})\Lambda^{2}}{2\pi^{2}}\sim 1.5$, which is fixed by fitting to QCD hadron observables (see also Eq.(22)). Thus, QCD indeed yields a significantly large soft-scale breaking as dynamical chiral symmetry breaking. Note that the near criticality condition, $\frac{N_{f}N_{c}(g_{s}+g_{d})\Lambda^{2}}{2\pi^{2}}\sim 1$, also implies the dynamics near an ultraviolet fixed point (in the chiral broken phase for $\frac{N_{f}N_{c}(g_{s}+g_{d})\Lambda^{2}}{2\pi^{2}}>1$), $\Lambda\frac{\partial[(g_{s}+g_{d})\Lambda^{2}]}{\partial\Lambda}\sim 0$, i.e., the almost quantum scale invariance.
Below we give two possible scenarios with some Beyond the Standard Model to achieve the CW-type chiral phase transition characterized by the near criticality soft-scale breaking in Eq.(12).
III.1 Walking dark QCD
In the MFA, the near criticality condition for the soft-scale breaking, $\frac{N_{f}N_{c}(g_{s}+g_{d})\Lambda^{2}}{2\pi^{2}}=1$, implies
$\frac{M\Lambda^{2}}{\langle-\bar{q}q\rangle/N_{f}}=\frac{4\pi^{2}}{N_{c}}\sim 12\times\left(\frac{3}{N_{c}}\right)\,,$ (13)
where we have used Eq.(II) with $\theta=m=0$. Two-flavor QCD gives $\frac{M\Lambda^{2}}{\langle-\bar{q}q\rangle/N_{f}}\sim 20$. For a fixed place of $S=\langle\bar{q}q\rangle$ in the thermodynamic potential, Eq.(13) requires a smaller dynamical (full) mass $M$ by about 40%, compared to the existing two-flavor QCD. Since generically $\langle-\bar{q}q\rangle\sim M^{3}$ (when evaluated at the renormalization scale $\mu=M$), one may read the condition in Eq.(13) as $\Lambda/M\gg 1$. This indicates a large scale hierarchy between the ultraviolet and infrared scales intrinsic to the underlying theory contrast to QCD: ordinary QCD yields almost the same order for them, $\Lambda\sim(2-3)M$. The soft-scale breaking is ensured by the nature of the almost scale-invariant gauge theory, and the CW-type scale anomaly is induced by nonperturbative running of the gauge coupling associated with the dynamical chiral symmetry breaking Leung:1985sn . One candidate theory to realize such a large scale gap is many flavor QCD, e.g., QCD with $N_{f}=8$ Aoki:2013xza ; LSD:2014nmn ; Hasenfratz:2014rna . In this scenario, the approximate quantum scale symmetry gets anomalous dominantly by the CW-type scale anomaly of the form $\sim M^{4}\ln M$. This characteristic feature has so far been applied to various phenomena, such as dynamical CW inflation scenarios Ishida:2019wkd ; Cacciapaglia:2023kat ; Zhang:2023acu ; Zhang:2024vpp ; Liu:2024xrh ; Zhang:2025kbu ; Cacciapaglia:2025xqd , models of dynamical electroweak symmetry breaking Matsuzaki:2012gd ; Matsuzaki:2012vc ; Matsuzaki:2012mk ; Matsuzaki:2013eva ; Matsuzaki:2015sya ; Kasai:2016ifi ; Hansen:2016fri ; Appelquist:2017wcg ; Appelquist:2017vyy ; Cata:2019edh ; Appelquist:2019lgk ; Brown:2019ipr ; Elander:2025fpk ; Faedo:2024zib ; Appelquist:2024koa ; Ingoldby:2023mtf ; Freeman:2023ket ; LSD:2023uzj ; Choi:2011fy ; Choi:2012kx ; Appelquist:2022mjb ; Appelquist:2022qgl ; Golterman:2020utm ; Miura:2018dsy , and phenomenological applications to QCD dilaton Kawaguchi:2020kce ; Crewther:2013vea ; Li:2016uzn ; Fujii:2025aip ; Sheng:2025jvn ; Zhang:2024iye ; Zhang:2024sju ; Bersini:2023uat ; Ma:2023ugl ; Shao:2022njr ; Cata:2019edh ; Li:2017udr ; Paeng:2017qvp ; Ma:2016gdd ; Ma:2016nki .
III.2 QCD with ALP
One may also consider a new particle ($\Phi$) coupled to $M$, like $\Phi^{2}M^{2}$, and allow $\Phi$ to develop the vacuum expectation value and generate an extra soft-scale breaking mass term, to cancel the dynamical chiral symmetry breaking at low-enough $T$. Suppose that this vacuum expectation value becomes nonzero in the chiral symmetric phase at high $T$, and almost vanishes at the vacuum $T=0$. This setup ensures that the extra-soft scale breaking effect is activated to trigger the cosmological CW-type phase transition at $T=T_{n}$, and below $T_{n}$, it freezes out to make the QCD-chiral broken phase normal, so that the present-day QCD hadron physics is intact. The decoupling of the extra-soft scale breaking can be realized, say, if the vacuum expectation value of $\Phi$ scales with $\mu_{5}$, the chiral chemical potential, so the $\Phi$ field acts like an ALP coupled to the axial charge, as has been discussed in Zhang:2025kbu . This decoupling is then encoded with the freezing out of the QCD sphaleron, which would happen in the epoch of the confinement-deconfinement phase transition, which takes place seemingly simultaneously with the chiral phase transition. As long as the QCD sphaleron is active in hot QCD plasma, the chiral chemical potential $\mu_{5}$ along with a local CP-odd domain is created Manton:1983nd ; Klinkhamer:1984di , so that the QCD vacuum characterized by the strong CP phase $\theta$ and its fluctuation (in the spatial-homogeneous direction) gets significantly sizable Kharzeev:2007tn ; Kharzeev:2007jp ; Fukushima:2008xe within the QCD time scale McLerran:1990de ; Moore:1997im ; Moore:1999fs ; Bodeker:1999gx : the sphaleron transition rate is not suppressed by the thermal effect in contrast to the instanton’s one Moore:1997im ; Moore:1999fs ; Bodeker:1999gx . The time fluctuation $\partial_{t}\theta(t)$, to be referred to as the chiral chemical potential, $\mu_{5}$ Kharzeev:2007tn ; Kharzeev:2007jp ; Fukushima:2008xe ; Andrianov:2012hq ; Andrianov:2012dj , will be significant as well when the non-conservation law of the $U(1)$ axial symmetry is addressed.
We assume $\Phi$ to be a pseudoscalar which couples to up and down quarks via the Yukawa coupling term,
$y\,\bar{q}i\gamma_{5}\Phi q\,.$ (14)
Then, the $\Phi$ coupling to the chiral chemical potential $\mu_{5}$ would also be induced through the axial covariance:
$\frac{1}{2}|D_{\mu}\Phi|^{2}\,,\qquad{\rm with}\qquad D_{\mu}\Phi=(\partial_{\mu}-i\mu_{5}\delta^{0}_{\mu})\Phi\,.$ (15)
Hence $\Phi$ gets the effective potential along with $\mu_{5}$
$V_{\mu_{5}}(S)=-\frac{1}{2}\mu_{5}^{2}\Phi^{2}\,.$ (16)
This is definitely a negative mass-squared term ensured by the gauge covariance for the axialvector field $A_{\mu}=\mu_{5}\delta_{\mu}^{0}$, which corresponds to the repulsive force from the vector interaction against the scalar probes charged under the corresponding gauge. Thus, combined with the quartic $\lambda\Phi^{4}$ term, $\Phi$ can get nonzero vacuum expectation value, $v_{\Phi}$.
The Yukawa coupling term in Eq.(14), when put in the current MFA-NJL framework, modifies the thermodynamic potential. It makes a shift of the pseudoscalar mean field $P$ terms by $(y\,\Phi)$. Keeping $\theta=0$, we then find that up to thermal correction parts, the vacuum contribution in Eq.(10) looks like modified as
$\Omega$ $\approx\frac{\widetilde{M^{2}}}{2(g_{s}+g_{d})}+\frac{g_{d}\widetilde{P}^{2}}{g_{s}^{2}-g_{d}^{2}}-\frac{N_{f}N_{c}}{4\pi^{2}}\left[\widetilde{M}^{2}\Lambda^{2}-\frac{(\widetilde{M}^{2})^{2}}{4}\right]\ln\frac{\Lambda^{2}}{\widetilde{M}^{2}}\,,$ (17)
where
$\widetilde{M}^{2}$ $\equiv M^{2}+\widetilde{P}^{2}=4(g_{s}+g_{d})^{2}S^{2}+\widetilde{P}^{2}\,,$
$\widetilde{P}^{2}$ $\equiv[y\Phi-2(g_{s}-g_{d})P]^{2}\,.$ (18)
In Eq.(17) we have take the chiral limit $m\to 0$ for simplicity. The $\mu_{5}$-induced mass term for $\Phi$ in Eq.(16) makes the total potential aligned to develop the vacuum expectation values for $\Phi$ and also $P$, both of which are proportional to $\mu_{5}$. Hence the total potential will be stabilized separately along $M$ and $\widetilde{P}$ directions. Since the Yukawa coupling in Eq.(14) is CP invariant, we assume $\langle P\rangle=0$ at $\theta=0$, as in ordinary QCD. Then, in the $M$ direction, the developed $\langle\Phi\rangle_{\mu_{5}}$, generates the extra positive mass-squared contribution to $M$,
$\sim\frac{N_{f}N_{c}}{8\pi^{2}}\langle\widetilde{P}\rangle_{\mu_{5}}^{2}\,,$ (19)
up to the logarithmic factor of ${\cal O}(1)$. Thus, the total $M^{2}$ term goes like
$\Omega_{M^{2}}$ $\sim\left[\frac{1}{2(g_{s}+g_{d})}-\frac{N_{f}N_{c}}{4\pi^{2}}\Lambda^{2}+\frac{N_{f}N_{c}}{8\pi^{2}}\langle y\,\Phi\rangle_{\mu_{5}}^{2}\right]M^{2}+\cdots\,,$
$\sim\frac{N_{f}N_{c}}{8\pi^{2}}\left[\langle y\,\Phi\rangle_{\mu_{5}}^{2}-(475\,{\rm MeV})^{2}\right]M^{2}+\cdots\,,$ (20)
where ellipses stand for thermal corrections terms at $T\gtrsim T_{c}$, and we have quoted values of the model parameters referenced in Eq.(22). As clearly seen from Eq.(20), the soft-scale breaking mass term can be canceled when $\langle y\,\Phi\rangle_{\mu_{5}}\sim 475$ MeV, so that the chiral phase transition almost goes like the CW-type supercooling. It is interesting to note also that this $\Phi$-ALP, within only or extending the present framework can be a candidate to solve the strong CP problem.
The ALP-Yukawa coupling to light quarks as in Eq.(14) has severely been constrained by the astrophysical observations and collider experiments, via the ALP decay to diphoton. See the public website at https://cajohare.github.io/AxionLimits or https://cajohare.github.io/AxionLimits/docs/ap.html. The diphoton coupling for the present $\Phi$-ALP is induced by mixing with the QCD-$\eta^{\prime}$ meson, like in the case of composite ALP models at around the QCD scale Wang:2024vlm . The mixing arises through the Yukawa coupling $y$ in Eq.(14), so the $\Phi-\gamma-\gamma$ coupling $g_{\Phi\gamma\gamma}$ can roughly be of ${\cal O}(y\cdot m_{\eta^{\prime}}^{2}/|m_{\Phi}^{2}-m_{\eta^{\prime}}^{2}|\cdot g_{\eta\gamma\gamma})$, where $g_{\eta\gamma\gamma}\sim 10^{-3}\,{\rm GeV}^{-1}$ (see, e.g., Wang:2024vlm ) based on a chiral effective model approach. Since $\mu_{5}={\cal O}(T_{c})={\cal O}(100-200\,{\rm MeV})$, the size of $y$ is desired to be $\lesssim{\cal O}(1)$ and cannot be extremely small, it is likely that the $\Phi$-ALP is phenomenologically as heavy as $\gtrsim 1$ GeV, otherwise it has already been ruled out. More detailed phenomenological and cosmological consequences would be worth investigating, to be pursued elsewhere.
IV Exploring CW-type chiral phase transition:: case with $\mu_{B}\neq 0$
At an extreme limit where $T\sim 0$ and finite $\mu_{B}$ (with $\mu_{5}=0$), the finite-density correction to $\Omega$ in Eq.(6) can be evaluated as
$-2N_{c}N_{f}\lim_{T\to 0}T\int\frac{d^{3}p}{(2\pi)^{3}}\left[\text{ln}\left(1+e^{-\frac{E_{0}+\frac{1}{3}\mu_{B}}{T}}\right)+\text{ln}\left(1+e^{-\frac{E_{0}-\frac{1}{3}\mu_{B}}{T}}\right)\right]$
$\approx\frac{N_{c}N_{f}}{24\pi^{2}}\left(-(\frac{1}{3}\mu_{B})\sqrt{(\frac{1}{3}\mu_{B})^{2}-M^{2}}\left(2\cdot(\frac{1}{3}\mu_{B})^{2}-5M^{2}\right)\right)$
$+\frac{N_{c}N_{f}}{24\pi^{2}}\left(-3M^{4}\ln{\frac{(\frac{1}{3}\mu_{B})+\sqrt{(\frac{1}{3}\mu_{B})^{2}-M^{2}}}{M}}\right)$
$\equiv\Omega_{\mu_{B}}^{(1)}+\Omega_{\mu_{B}}^{(2)}\,.$ (21)
The first term defined as $\Omega_{\mu_{B}}^{(1)}$, up to a term constant in $M$, goes like $\sim\frac{N_{c}N_{f}}{12\pi^{2}}(\frac{\mu_{B}}{3})^{2}M^{2}$ for a large $\mu_{B}\gg M$, which is the Debye mass screening. The second term $\Omega_{\mu_{B}}^{(2)}\sim\frac{-N_{c}N_{f}}{16\pi^{2}}M^{4}\ln\frac{\left[(\frac{\mu_{B}}{3})+\sqrt{(\frac{\mu_{B}}{3})^{2}-M^{2}}\right]^{2}}{M^{2}}$ takes a scale anomaly form of the CW type and, in an extreme limit $\mu_{B}\gg M$, precisely cancels the original one at $\mu_{B}=0$ in Eq.(9). A potential barrier is generated at moderately large $\mu_{B}$ by creation of the Fermi surface, acting as a repulsive force, which is dictated by $\Omega_{\mu_{B}}^{(1)}$. This triggers the first order phase transition at $T=0$.
At finite $T$, the cancellation for the original scale anomaly form of the CW-type becomes incomplete, hence it will survive during the thermal phase transition, as long as the Debye mass screening by $\mu_{B}$ efficiently works so as to make the dynamical chiral symmetry breaking milder. Thus, a large enough $\mu_{B}$ can control the size of the soft-scale breaking, and at finite $T$, can pull the type of the first order phase transition back to the CW-type.
Figure 1 demonstrates how the CW-type potential can be realized by finite $\mu_{B}$ at low enough $T$. The appropriate size of $\mu_{B}$ would thus be $\sim 1100$ MeV, which is much greater than the typical QCD chemical potential size of ${\cal O}(300)$ MeV. Requiting this size of $\mu_{B}$ is also consistent with the analysis in the literature  Shao:2024dxt ; Shao:2024ygm , to achieve sizable nano hertz GW production from quark-meson models. Note that in Fig. 1, the soft-scale breaking effect $\propto S^{2}$ contribute negatively at around $S\sim 0$. Here, to fix the model parameters, we have referred to empirical hadron observables in the isospin symmetric limit at the vacuum Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs : the pion mass $m_{\pi}=140.2\,\mathrm{MeV}$, the pion decay constant $f_{\pi}=92.6\,\mathrm{MeV}$, and quark condensate $\langle\bar{u}u\rangle=\langle\bar{d}d\rangle=(-241.5\,\mathrm{MeV})^{3}$. Then the model parameters can be determined as Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs ,
$m\simeq 6\,\mathrm{MeV},\qquad\Lambda\simeq 590\,\mathrm{MeV}\,,$
$g_{s}=2(1-c)G_{0},\qquad g_{d}=2cG_{0}\,,\qquad{\rm with}\qquad c\simeq 0.2\,,\qquad G_{0}\Lambda^{2}\simeq 2.435\,.$ (22)
In Fig. 2 we show the thermal evolution of the CW-type chiral phase transition with $\mu_{B}=1100$ MeV fixed. The CW supercooling can indeed be achieved in this setup. The stochastic GW background can then be produced and the predicted spectra could be similar to those discussed in Shao:2024dxt ; Shao:2024ygm , where a two-flavor quark meson model with a large $\mu_{B}$ such as $\mu_{B}\sim 1$ GeV has been shown to generate nano hertz GW signals, without clarifying the limit to the CW-type chiral phase transition. As inspired also in the literature Shao:2024dxt ; Shao:2024ygm , the Affleck-Dine baryogenesis Affleck:1984fy ; Linde:1985gh ; Dine:2003ax could produce such a large $\mu_{B}$ in the QCD phase transition epoch.
Figure 1: The variation of the thermodynamic potential at low $T$ ($T=20$ MeV) with respect to $\mu_{B}$, in unit of ${\rm MeV}^{4}$, as a function of $S=\frac{M-m}{2(g_{s}+g_{d})}$ at $\theta=0$ in unit of ${\rm MeV}^{3}$, in the MFA-NJL description monitoring two-flavor QCD at physical point (for details, see the text).
Figure 2: The thermal chiral phase transition of the CW type including a soft-scale breaking mass term induced from the Debye mass at $\mu_{B}=1100$ MeV and dynamical chiral symmetry breaking. The model parameters have been setup at physical point as in Fig. 1.
V Conclusion
In conclusion, we have clarified that the chiral phase transition in QCD-like theories can be Coleman-Weinberg type supercooling. This possibility paves a way to a new search window for new physics to be probed by gravitational wave and primordial black hole productions. We have discussed this possibility by monitoring ordinary QCD setup in a view of a two-flavor Nambu-Jona-Lasinio model in hot and/or dense medium, at the mean-field approximation level. It turned out that the identification essentially follows the scale violation classification: soft-scale breaking and quantum scale anomaly of the CW-type. The Coleman-Weinberg type chiral phase transition is possible to realize when ordinary QCD is coupled to or replaced by Beyond the Standard Model in a nontrivial way. One candidate scenario includes prediction of an axionlike particle, which is likely to have a rich cosmology and phenomenological consequences at around the QCD scale. More detailed study would be worth performing in the future.
Arguments similar to the present Nambu-Jona-Lasinio’s description can also be applied to bosonic models, e.g., quark meson models and linear sigma models, or more generic boson-fermion system. The key ingredient is at any rate identifying the CW-type scale anomaly and the size of the soft-scale breaking. Such a more general investigation would also be noteworthy.
Acknowledgments
This work was supported in part by the National Science Foundation of China (NSFC) under Grant No.11747308, 11975108, 12047569, and the Seeds Funding of Jilin University (S.M.). The work by M.K. is supported by RFIS-NSFC under Grant No. W2433019. The work of A.T. was partially supported by JSPS KAKENHI Grant Numbers 20K14479, 22K03539, 22H05112, and 22H05111, and MEXT as “Program for Promoting Researches on the Supercomputer Fugaku” (Simulation for basic science: approaching the new quantum era; Grant Number JPMXP1020230411, and Search for physics beyond the standard model using large-scale lattice QCD simulation and development of AI technology toward next-generation lattice QCD; Grant Number JPMXP1020230409).